# MSACL: Multi-Step Actor-Critic Learning with Lyapunov Certificates for Exponentially Stabilizing Control

### v2

@@ Abstract
For safety-critical applications, model-free reinforcement learning (RL) faces numerous challenges, particularly the difficulty of establishing verifiable stability guarantees while maintaining high exploration efficiency. To address these challenges, we present Multi-Step Actor-Critic Learning with Lyapunov Certificates (MSACL), a novel approach that seamlessly integrates exponential stability with maximum entropy reinforcement learning (MERL). In contrast to existing methods that rely on complex reward engineering and single-step constraints, MSACL utilizes intuitive rewards and multi-step data for actor-critic learning. Specifically, we first introduce Exponential Stability Labels (ESLs) to categorize samples and propose a $\lambda$-weighted aggregation mechanism to learn Lyapunov certificates. Leveraging these certificates, we then develop a stability-aware advantage function to guide policy optimization, thereby ensuring rapid Lyapunov descent and robust state convergence. We evaluate MSACL across six benchmarks, comprising four stabilization and two high-dimensional tracking tasks. Experimental results demonstrate its consistent superiority over both standard RL baselines and state-of-the-art Lyapunov-based RL algorithms. Beyond rapid convergence, MSACL exhibits significant robustness against environmental uncertainties and remarkable generalization to unseen reference signals. The source code and benchmarking environments are available at https://github.com/YuanZhe-Xing/MSACL.
@@ I Introduction
Reinforcement learning (RL) has emerged as a transformative paradigm for the control of complex, nonlinear dynamical systems [1, 2], demonstrating remarkable empirical success across high-dimensional tasks without requiring explicit analytical models [3, 4]. Despite these advancements, the deployment of RL in safety-critical applications, such as autonomous driving and human-robot interaction, remains fundamentally constrained by a lack of formal reliability guarantees [5, 6].
While alternative frameworks like constrained Markov Decision Processes (cMDPs) [7, 8] and Hamilton-Jacobi reachability analysis [9, 10] offer potential solutions, certificate-based methods constitute a fundamental framework for verifying and stabilizing learning agents. By leveraging rigorous control-theoretic tools, such as Lyapunov certificates for establishing asymptotic stability [11, 12, 14, 13] and barrier certificates for ensuring set invariance [15, 16], these methods provide concise, data-driven proofs of system behavior. The synthesis of stability certificates has undergone a paradigm shift from computationally intensive analytical methods toward scalable neural architectures. Traditional approaches, such as sum-of-squares (SOS) optimization [17, 18, 19] and algebraic synthesis[20], are often restricted to polynomial dynamics. To bridge the gap between black-box learning and verifiable control, the field has evolved from probabilistic methods, notably the use of Gaussian Processes for safe exploration [21], to formal frameworks for synthesizing neural certificates that estimate regions of attraction (ROA) [22, 23]. Building on this verification-centric perspective, recent research has advanced toward the joint synthesis of neural certificates and policies. For instance, Chang et al. [24, 25] and Quan et al. [26, 27, 28] demonstrated that neural certificates can function as critics during training, directly guiding policy optimization to satisfy formal stability constraints.
@@
In contrast to the stability-by-design paradigms presented above, standard model-free RL algorithms, such as DDPG [29], PPO [31], TD3 [32], SAC [33], and SAC’s distributional variant DSAC [34, 35], lack intrinsic stability guarantees, thus facing fundamental limitations in precision control tasks. These methods treat stability as an incidental byproduct of reward maximization rather than a structural property. Consequently, employing general RL algorithms for safety-critical tasks necessitates extensive reward engineering to achieve stable behaviors [1, 2]. To bridge this gap, Lyapunov-based RL algorithms have been developed to integrate stability constraints into the learning process. For instance, Policy Optimization with Lyapunov Certificates (POLYC) [25] incorporates Lyapunov certificates within the PPO framework, while Lyapunov-based Actor-Critic (LAC) [36] introduces an off-policy approach for learning stability-constrained critics. By augmenting standard RL architectures with learned Lyapunov certificates, these methods provide structural stability guarantees for nonlinear control systems. However, several critical limitations persist. First, these approaches suffer from low sample efficiency under on-policy training paradigms. Second, they often fail to account for control costs during policy optimization. More critically, most existing methods focus solely on asymptotic stability, which lacks the rapid stabilization characteristics necessary for high-performance response in dynamic control scenarios.
To overcome these limitations, we propose Multi-Step Actor-Critic Learning with Lyapunov Certificates (MSACL), a Lyapunov-based off-policy RL algorithm that enhances sample efficiency [37, 38] while optimizing a comprehensive objective that balances state regulation and control cost. By integrating exponential stability constraints into the maximum entropy reinforcement learning (MERL) architecture, MSACL ensures rapid stabilization and enhanced robustness. A key feature of our approach is the multi-step learning mechanism, which captures long-horizon dynamics to establish connections between temporal difference learning and formal stability verification. By mitigating discretization errors and stochastic noise associated with single-step transitions, this mechanism enables the synthesis of policies that are both high-performing and mathematically certifiable.
In summary, this article presents a Lyapunov-based RL approach for model-free exponentially stabilizing control. The main contributions are summarized as follows:
@@
    Exponential Stability via Multi-Step Learning. We propose the MSACL algorithm, which integrates exponential stability with Lyapunov certificate learning via a multi-step mechanism. This algorithm enforces exponential convergence rates, providing stability-guaranteed performance for complex nonlinear control, surpassing conventional asymptotic methods.
@@
    Sample-Efficient Off-Policy Framework. We develop a sample-efficient off-policy framework that leverages the maximum entropy paradigm and importance sampling. This framework effectively manages distribution shifts and interaction data reuse, thus enhancing training stability compared to existing on-policy algorithms.
@@
    Systematic Evaluation and Robustness Analysis. We establish an evaluation platform across six benchmarks comprising stabilization and high-dimensional tracking tasks. Experiment results demonstrate superior performance, robustness and generalization of MSACL, while sensitivity analyses elucidate the multi-step bias-variance trade-off to guide diverse control tasks.
@@ II Preliminaries
This section establishes the theoretical foundations for MSACL. We first define the stabilization and tracking problems for model-free discrete-time systems, then introduce Lyapunov-based exponential stability criteria, and finally review the MERL framework.
@@ II-A Problem Statement
In this article, we study the stabilization problem for general discrete-time deterministic control systems:
@@
where $\mathbf{x}_{t}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ denotes the state at time $t$, $\mathbf{u}_{t}\in\mathcal{U}\subseteq\mathbb{R}^{m}$ denotes the control input, $\mathcal{X}$ and $\mathcal{U}$ represent the sets of admissible states and control inputs, respectively, and $f:\mathcal{X}\times\mathcal{U}\to\mathcal{X}$ is the flow map. Given that $f$ is unknown, we adopt a model-free control paradigm. Let $\mathcal{X}_{0}\subseteq\mathcal{X}$ denote the set of initial states. For any $\mathbf{x}_{0}\in\mathcal{X}_{0}$, let $\mathbf{x}(t,\mathbf{x}_{0})$ denote the state at time $t$ along the trajectory starting from $\mathbf{x}_{0}$. For notational simplicity, $\mathbf{x}_{t}$ is used to denote $\mathbf{x}(t,\mathbf{x}_{0})$ whenever the context is unambiguous. It is worth noting that the discrete-time case presented in this article can be similarly extended to the continuous-time case; interested readers may refer to [11] and [25] for details.
In the context of control systems, our objective is to find a state feedback policy $\pi:\mathcal{X}\to\mathcal{U}$ such that the control input $\mathbf{u}_{t}=\pi(\mathbf{x}_{t})$ renders the closed-loop system $\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\pi(\mathbf{x}_{t}))$ to satisfy exponential stability. For clarity, we consider the scenario where the state $\mathbf{x}_{t}$ is fully observable, and the policy $\pi$ is stationary, i.e., it depends only on the state and not explicitly on time. Specifically, the problem studied in this article is formulated as:
@@
(Exponential Stabilization) For the control system (1), design a stationary policy $\pi$ such that for any initial state $\mathbf{x}_{0}\in\mathcal{X}_{0}$, the resulting closed-loop trajectory $\mathbf{x}(t,\mathbf{x}_{0})$ exponentially converges to the equilibrium state $\mathbf{x}_{g}$.
Without loss of generality, let $\mathbf{x}_{g}$ denote the origin, i.e., $\mathbf{x}_{g}=\mathbf{0}$. Furthermore, the tracking problem is inherently encompassed within the stabilization framework, as tracking a time-varying reference signal $\mathbf{x}_{t}^{\text{ref}}$ reduces to stabilizing the error $\mathbf{e}_{t}=\mathbf{x}_{t}-\mathbf{x}_{t}^{\text{ref}}$ at zero.
@@ II-B Exponential Stability and Lyapunov Certificates
To guarantee that the closed-loop system is exponentially stable with respect to $\mathbf{x}_{g}$, we introduce the following lemma:
@@ II-C Standard RL and Maximum Entropy RL
Furthermore, we employ stochastic policies in this article to facilitate exploration within the maximum entropy RL (MERL) framework. A stochastic policy is defined as a mapping $\pi:\mathcal{X}\to\mathcal{P}(\mathcal{U})$, where $\mathcal{P}(\mathcal{U})$ denotes the set of probability distributions over $\mathcal{U}$, and $\mathbf{u}_{t}\sim\pi(\cdot|\mathbf{x}_{t})$ is sampled from the probability distribution. We denote the state and state-action marginals induced by $\pi$ as $\rho_{\pi}(\mathbf{x}_{t})$ and $\rho_{\pi}(\mathbf{x}_{t},\mathbf{u}_{t})$, respectively. In contrast to standard RL, which maximizes the expected cumulative discounted reward, MERL optimizes an entropy-augmented objective [33, 34, 35]:
@@ III-A Data Collection
In contrast to conventional Lyapunov-based RL algorithms [25, 36], which primarily target asymptotic stability through single-step transitions, our proposed MSACL enforces the stricter exponential stability requirement. Guided by the multi-step relationship (4), establishing exponential stability necessitates leveraging multi-step sequence samples from state trajectories. This sequential data structure provides a sufficient temporal horizon to validate the contraction properties of Lyapunov certificates.
@@ III-B Loss Function Design of Lyapunov Network
On the other hand, learning Lyapunov certificates directly from off-policy data may not be accurate, as empirical sequences in the early stages of training often fail to satisfy the exponential decay condition (5), leading to unreliable certificate estimation. To establish a rigorous correspondence between empirical data and theoretical stability requirements, we propose the Exponential Stability Label (ESL) mechanism. Specifically, we introduce positive and negative samples to bridge the gap between theory and practice:
@@ IV-B Loss Function Design of Policy Network
This subsection details the policy network optimization, which seamlessly integrates the maximum entropy objective with a novel Lyapunov-based stability guidance mechanism.
@@ IV-C Unified Actor-Critic Optimization Framework
Algorithm 1 incorporates two implementation refinements. First, for the soft Q-loss (14), we replace $Q_{\theta}$ with $\min\{Q_{\theta_{1}},Q_{\theta_{2}}\}$ to mitigate overestimation bias. Second, we implement delayed updates for both policy $\pi_{\phi}$ and temperature $\alpha$, performing $d$ consecutive updates within each delayed cycle to ensure sufficient policy improvement. These techniques, adapted from the cleanrl framework [42], significantly improve training stability and performance.
@@ V Experiments
This section evaluates the MSACL algorithm across six benchmarking environments, covering stabilization and high-dimensional tracking tasks. We compare our approach against model-free and Lyapunov-based RL baselines using performance and stability metrics. Further analysis includes Lyapunov certificate visualization, robustness assessment under parametric uncertainty and stochastic noise, and sensitivity studies on the multi-step horizon $n$.
@@ V-A Benchmarking Environments
This subsection presents the benchmarking environments for evaluating MSACL and baseline algorithms. We categorize environments by control objective: stabilization tasks include Controlled VanderPol (VanderPol)[43], Pendulum[44], Planar DuctedFan (DuctedFan)[45], and Two-link planar robot (Two-link)[46]; tracking tasks comprise SingleCarTracking[47] and QuadrotorTracking[48], as illustrated in Figure 3.
For stabilization tasks, the goal is to drive the state to the origin $\mathbf{x}_{g}=\mathbf{0}\in\mathbb{R}^{d}$; for tracking tasks, the system follows a time-varying reference signal $\mathbf{x}_{t}^{\text{ref}}$. All tasks use a unified reward structure:
@@ V-B Baseline Algorithms
We evaluate MSACL against classic model-free RL baselines (SAC [33], PPO [31]) and state-of-the-art Lyapunov-based algorithms (LAC [36], POLYC [25]). SAC and PPO serve as performance baselines for stabilization and tracking tasks, while LAC and POLYC provide stability guarantees through Lyapunov constraints. For LAC, the standard reward is replaced with a quadratic cost function defined as:
@@
where $\mathbf{Q_{c}}$ is positive definite. Note that in contrast to this state-only penalty formulation, our reward $r_{t}$ accounts for both state deviations and control effort. By comparing against these baselines, we aim to demonstrate the advantages of MSACL in terms of performance and stability metrics.
For fair comparison, all algorithms are implemented in the GOPS framework [51]. Actor and critic networks utilize MLPs with two 256-unit hidden layers and ReLU activations. All actors employ stochastic diagonal Gaussian policies. Critic architectures vary: LAC uses only a Lyapunov network, while POLYC incorporates both value and Lyapunov functions. SAC and LAC also implement clipped double Q-learning and delayed policy updates according to Remark 3. Optimization is performed via the Adam optimizer [30], with core hyperparameters listed in Table I.

### v3

@@ Abstract
For stabilizing control tasks, model-free reinforcement learning (RL) approaches face numerous challenges, particularly regarding the issues of effectiveness and efficiency in complex high-dimensional environments with limited training data. To address these challenges, we propose Multi-Step Actor-Critic Learning with Lyapunov Certificates (MSACL), a novel approach that integrates exponential stability into off-policy maximum entropy reinforcement learning (MERL). In contrast to existing RL-based approaches that depend on elaborate reward engineering and single-step constraints, MSACL adopts intuitive reward design and exploits multi-step samples to enable exploratory actor-critic learning. Specifically, we first introduce Exponential Stability Labels (ESLs) to categorize training samples and propose a $\lambda$-weighted aggregation mechanism to learn Lyapunov certificates. Based on these certificates, we further design a stability-aware advantage function to guide policy optimization, thereby promoting rapid Lyapunov descent and robust state convergence. We evaluate MSACL across six benchmarks, comprising four stabilizing and two high-dimensional tracking tasks. Experimental results demonstrate its consistent performance improvements over both standard RL baselines and state-of-the-art Lyapunov-based RL algorithms. Beyond rapid convergence, MSACL exhibits robustness against environmental uncertainties and generalization to unseen reference signals. The source code and benchmarking environments are available at https://github.com/YuanZhe-Xing/MSACL.
@@ I Introduction
Stabilizing control is a central requirement in real-world control systems, especially in safety-critical applications such as autonomous driving [1], aerospace systems [2], and robotics [3]. In these settings, a controller must not only accomplish the desired task but also guarantee stable system behavior under uncertainty and disturbances [4, 5]. As control systems become increasingly complex and are deployed in dynamic, partially unknown environments, developing control strategies that can simultaneously ensure stability and scalability has become a fundamental challenge [6]. Traditionally, control strategies have relied heavily on rigorous mathematical frameworks and precise analytical models[7, 8], such as constrained Markov Decision Processes (cMDPs) [9], Hamilton-Jacobi reachability analysis [10] and certificate-based methods utilizing strict control-theoretic tools (e.g., sum-of-squares optimization [11, 12] and algebraic synthesis [13]). However, all these traditional approaches face severe scalability challenges when applied to complex nonlinear systems with unknown dynamics.
@@
To address the complexities of real-world dynamical systems without requiring explicit analytical models, RL has emerged as a highly effective paradigm for optimal control [14], demonstrating considerable empirical success across various control tasks [15, 16]. Nevertheless, standard model-free RL algorithms, such as DDPG [17], PPO [18], and SAC [19], lack intrinsic stability awareness. They treat stability as an incidental byproduct of reward maximization, necessitating extensive reward engineering to achieve stable or safe behaviors. To bridge this gap, Lyapunov-based RL algorithms have been developed to integrate stability constraints into the learning process. For instance, Policy Optimization with Lyapunov Certificates (POLYC) [20] incorporates Lyapunov certificates within the PPO framework, while Lyapunov-based Actor-Critic (LAC) [21] introduces an approach for learning stability-constrained critics. By augmenting standard RL architectures with learned neural Lyapunov certificates, these methods empirically facilitate structural stability for nonlinear control systems.
However, despite these advances, existing Lyapunov-based RL methods often struggle with both effectiveness and efficiency in complex, high-dimensional stabilizing control tasks (e.g., Quadrotor tracking). The primary bottleneck for effectiveness lies in their inability to learn sufficiently accurate Lyapunov certificates. First, existing methods rely on single-step constraints, providing limited long-horizon information and leading to myopic policy improvement. Second, on-policy methods like POLYC depend exclusively on newly collected data, restricting the diversity of behaviors needed to accurately shape the Lyapunov landscape. Consequently, the learned certificates fail to provide reliable structural guidance, causing performance degradation or divergence.
To overcome the effectiveness issue, introducing a multi-step off-policy approach is essential, as it captures long-horizon trajectory evolution. Meanwhile, the multi-step off-policy approach naturally resolves the secondary issue: it boosts efficiency by deeply reusing historical experiences, overcoming the low sample efficiency inherent in previous on-policy methods. Inspired by the idea of using multi-step off-policy, we in this paper attempt to learn a high-fidelity Lyapunov certificate that can reliably guide policy updates. To this end, we propose Multi-Step Actor-Critic Learning with Lyapunov Certificates (MSACL). First, MSACL exploits multi-step samples to learn Lyapunov certificates, enabling the certificates to capture long-horizon system dynamics rather than relying only on local single-step transitions. Second, MSACL is built upon an off-policy maximum entropy reinforcement learning (MERL) framework, which not only improves sample efficiency through data reuse but also promotes broader exploration, thereby providing more diverse training data for characterizing system behaviors. While parameterizing Lyapunov certificates via neural networks inherently lacks the strict formal mathematical guarantees of traditional analytical methods, the proposed multi-step and off-policy learning paradigm yields a more reliable approximation empirically. This enhanced Lyapunov certificate, in turn, provides structurally grounded, stability-aware advantage estimates to guide policy optimization toward rapid and robust stabilization.
The proposed MSACL approach makes the following contributions to safe model-free control:
@@
    Multi-Step Learning with Lyapunov Guidance. We introduce a multi-step learning mechanism to learn high-precision Lyapunov certificates that better capture long-horizon system dynamics and provide more effective guidance for policy updates. Built upon this mechanism, MSACL further incorporates exponential stability through Exponential Stability Labels (ESLs) and a $\lambda$-weighted aggregation scheme, yielding stronger stability supervision and faster convergence than conventional single-step methods.
@@
    Sample-Efficient Off-Policy Framework. We develop MSACL within an off-policy MERL framework, moving beyond the data inefficiency of existing on-policy Lyapunov-based methods. By improving data reuse and encouraging broader exploration, the proposed framework captures more diverse system behaviors and provides richer training samples for Lyapunov certificate learning. In addition, importance sampling is incorporated to mitigate distribution shifts, further enhancing sample efficiency and training stability.
@@
    Systematic Evaluation and Robustness Validation. We establish a comprehensive evaluation platform across six benchmarks, including high-dimensional stabilizing control tasks. Experimental results demonstrate MSACL’s consistent empirical superiority over both standard RL and state-of-the-art Lyapunov-based baselines. Furthermore, MSACL exhibits robustness against environmental noise and generalization to unseen reference signals.
@@ II Preliminaries
This section establishes the theoretical foundations for MSACL. We first define the exponentially stabilizing control problems for model-free discrete-time systems, then introduce Lyapunov-based exponential stability criteria, and finally review the MERL framework.
@@ II-A Problem Statement
In this article, we study the stabilizing problem for general discrete-time deterministic control systems:
@@
where $\mathbf{x}_{t}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ denotes the state at time $t$, $\mathbf{u}_{t}\in\mathcal{U}\subseteq\mathbb{R}^{m}$ denotes the control input, $\mathcal{X}$ and $\mathcal{U}$ represent the sets of admissible states and control inputs respectively, and $f:\mathcal{X}\times\mathcal{U}\to\mathcal{X}$ is the flow map. Given that $f$ is unknown, we adopt a model-free control paradigm. Let $\mathcal{X}_{0}\subseteq\mathcal{X}$ denote the set of initial states. For any $\mathbf{x}_{0}\in\mathcal{X}_{0}$, let $\mathbf{x}(t,\mathbf{x}_{0})$ denote the state at time $t$ along the trajectory starting from $\mathbf{x}_{0}$. For notational simplicity, $\mathbf{x}_{t}$ is used to denote $\mathbf{x}(t,\mathbf{x}_{0})$ whenever the context is unambiguous. It is worth noting that the discrete-time case presented in this article can be similarly extended to the continuous-time case; interested readers may refer to [22] and [20] for details.
For the control system (1), our objective is to find a state feedback policy $\pi:\mathcal{X}\to\mathcal{U}$ such that the control input $\mathbf{u}_{t}=\pi(\mathbf{x}_{t})$ renders the closed-loop system $\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\pi(\mathbf{x}_{t}))$ to satisfy exponential stability. For clarity, we consider the scenario where the state $\mathbf{x}_{t}$ is fully observable, and the policy $\pi$ is stationary, i.e., it depends only on the state and not explicitly on time. Specifically, the problem studied in this article is formulated as:
@@
(Exponential stabilizing control) For the control system (1), design a policy $\pi$ such that for any initial state $\mathbf{x}_{0}\in\mathcal{X}_{0}$, the resulting closed-loop trajectory $\mathbf{x}(t,\mathbf{x}_{0})$ exponentially converges to the equilibrium state $\mathbf{x}_{g}$.
Without loss of generality, let $\mathbf{x}_{g}$ denote the origin, i.e., $\mathbf{x}_{g}=\mathbf{0}$. Furthermore, the tracking problem is inherently encompassed within the stabilizing framework, as tracking a time-varying reference signal $\mathbf{x}_{t}^{\text{ref}}$ reduces to stabilizing the error $\mathbf{e}_{t}=\mathbf{x}_{t}-\mathbf{x}_{t}^{\text{ref}}$ at zero.
@@ II-B Exponential Stability and Lyapunov Certificates
To facilitate that the closed-loop system is exponentially stable with respect to $\mathbf{x}_{g}$, we introduce the following lemma:
@@ II-C Standard RL and Maximum Entropy RL
Furthermore, we employ stochastic policies in this article to facilitate exploration within the MERL framework. A stochastic policy is defined as a mapping $\pi:\mathcal{X}\to\mathcal{P}(\mathcal{U})$, where $\mathcal{P}(\mathcal{U})$ denotes the set of probability distributions over $\mathcal{U}$, and $\mathbf{u}_{t}\sim\pi(\cdot|\mathbf{x}_{t})$ is sampled from the probability distribution. We denote the state and state-action marginals induced by $\pi$ as $\rho_{\pi}(\mathbf{x}_{t})$ and $\rho_{\pi}(\mathbf{x}_{t},\mathbf{u}_{t})$, respectively. In contrast to standard RL, which maximizes the expected cumulative discounted reward, MERL optimizes an entropy-augmented objective [19, 24, 25]:
@@ III-A Data Collection
In contrast to conventional Lyapunov-based RL algorithms [20, 21], which primarily target asymptotic stability through single-step transitions, our proposed MSACL enforces the stricter exponential stability requirements. Guided by the multi-step relationship (4), establishing exponential stability necessitates leveraging multi-step sequence samples from state trajectories. This sequential data structure provides a sufficient temporal horizon to effectively learn the contraction properties of Lyapunov certificates.
@@ III-B Loss Function Design of Lyapunov Network
On the other hand, learning Lyapunov certificates directly from off-policy data may not be accurate, as empirical sequences in the early stages of training often fail to satisfy the exponential decay condition (5), leading to unreliable certificate estimation. To establish a strong empirical alignment between empirical data and theoretical stability requirements, we propose the Exponential Stability Label (ESL) mechanism. Specifically, we introduce positive and negative samples to bridge the gap between theory and practice:
@@ IV-B Loss Function Design of Policy Network
This subsection details the policy network optimization, which integrates the maximum entropy objective with a novel Lyapunov-based stability guidance mechanism.
@@ IV-C Unified Actor-Critic Optimization Framework
Algorithm 1 incorporates two implementation refinements. First, for the soft Q-loss (14), we replace $Q_{\theta}$ with $\min\{Q_{\theta_{1}},Q_{\theta_{2}}\}$ to mitigate overestimation bias. Second, we implement delayed updates for both policy $\pi_{\phi}$ and temperature $\alpha$, performing $d$ consecutive updates within each delayed cycle to ensure sufficient policy improvement. These techniques, adapted from the cleanrl framework [31], improve training stability and performance.
@@ V Experiments
This section evaluates the MSACL algorithm across six benchmarking environments, covering stabilizing and high-dimensional tracking tasks. We compare our approach against model-free and Lyapunov-based RL baselines using performance and stability metrics. Further analysis includes Lyapunov certificate visualization, robustness assessment under parametric uncertainty and stochastic noise, and sensitivity studies on the multi-step horizon $n$.
@@ V-A Benchmarking Environments
This subsection presents the benchmarking environments for evaluating MSACL and baseline algorithms. We categorize environments by control objective: stabilizing tasks include Controlled VanderPol (VanderPol) [32], Pendulum [33], Planar DuctedFan (DuctedFan) [34], and Two-link planar robot (Two-link) [35]; tracking tasks comprise SingleCarTracking [36] and QuadrotorTracking [37], as illustrated in Figure 3.
For stabilizing tasks, the goal is to drive the state to the origin $\mathbf{x}_{g}=\mathbf{0}\in\mathbb{R}^{d}$; for tracking tasks, the system follows a time-varying reference signal $\mathbf{x}_{t}^{\text{ref}}$. All tasks use a unified reward structure:
@@ V-B Baseline Algorithms
We evaluate MSACL against classic model-free RL baselines (SAC [19], PPO [18]) and state-of-the-art Lyapunov-based algorithms (LAC [21], POLYC [20]). SAC and PPO serve as performance baselines for stabilizing and tracking tasks, while LAC and POLYC facilitate stability through Lyapunov constraints. For LAC, the standard reward is replaced with a quadratic cost function defined as:
@@
where $\mathbf{Q_{c}}$ is positive definite. Note that in contrast to this state-only penalty formulation, our reward $r_{t}$ accounts for both state deviations and control effort. By comparing against these baselines, we aim to demonstrate the advantages of MSACL in terms of performance and stability metrics introduced in the next subsection.
For fair comparison, all algorithms are implemented in the GOPS framework [40]. Actor and critic networks utilize MLPs with two 256-unit hidden layers and ReLU activations. All actors employ stochastic diagonal Gaussian policies. Critic architectures vary: LAC uses only a Lyapunov network, while POLYC incorporates both value and Lyapunov networks. SAC and LAC also implement clipped double Q-learning and delayed policy updates according to Remark 3. Optimization is performed via the Adam optimizer [41], with core hyperparameters listed in Table I.